# Graphical Abstract Highlights

### v1

@@ Abstract
West Nile virus (WNV) is a climate-sensitive mosquito-borne arbovirus circulating between mosquitoes of the genus Culex and birds, with a potential spillover to humans and other mammals. Recent trends in climatic change, characterized by early and/or prolonged summer seasons, increased temperatures, and above-average rainfall, probably facilitated the spread of WNV in Europe including Germany. In this work, we formulate a spatial WNV model consisting of a system of parabolic partial differential equations (PDEs), using the concept of diffusion and advection in combination with temperature-dependent parameters, i.e. mosquito biting rate, extrinsic incubation, and mortality rate. Diffusion represents the random movement of both mosquitoes and hosts across space, while advection captures the directed movement of migratory birds. The model is first studied mathematically, and we show that it has non-negative, unique, and bounded solutions in time and space. Numerical simulations of the PDE model are performed using temperature data for Germany (2019 - 2024). Results obtained from the simulation showed a high agreement with the reported WNV cases among birds and equids in Germany. The observed spreading patterns from the year 2018 to 2022 and the year 2024 were mainly driven by temperature in combination with diffusion processes of hosts and vectors. Only during the year 2023, the additional inclusion of advection for migratory birds was important to correctly predict new hotspots in new locations in Germany.
@@ 1 Introduction
West Nile virus (WNV) is an arbovirus of the Japanese encephalitis antigenic complex of the Flaviviridae family. It was first detected in a febrile patient from the West Nile district of Northern Uganda in 1937 (Smithburn et al., 1940), but is now considered the most geographically widespread arbovirus in the world (Ciota, 2017). The virus circulates in an enzootic cycle between Culex mosquitoes and birds, with regular spillovers to humans, equids, and other mammals (Calisher et al., 1989; Hubálek and Halouzka, 1999; Kramer et al., 2008). Birds are considered amplifying hosts as they develop sufficient levels of viremia, allowing for the infection of mosquitoes, while humans, equids, and other mammals are considered dead-end hosts (Rappole and Hubálek, 2003). Most humans infected with WNV are asymptomatic, but few may develop fever, headache, rash, among others (Petersen et al., 2013), including the risk of a fatal outcome (Campbell et al., 2002). Infected equids often exhibit ataxia, lethargy, and weakness as signs of the infection (Salazar et al., 2004). Multiple theories have been discussed regarding the introduction of new WNV cases in a previously unaffected region. There is evidence that migratory birds play an important role in the spread of the pathogen (Rappole and Hubálek, 2003). For example, birds captured during their seasonal migration have been tested positive for WNV in Israel (Ernek et al., 1977; Malkinson et al., 2002; Nir et al., 1967; Watson et al., 1972). Migratory birds likely spread the virus to stopover sites in Europe as they travel from the south to the north(Hannoun et al., 1972; Hubálek and Halouzka, 1999). In contrast, considering the speed and duration of flights of migratory birds in combination with the duration of viremia, other researchers argued that the short-distance movement of non-migratory birds better explains the observed spread pattern of WNV (Rappole and Hubálek, 2003).
@@
Several mathematical models for WNV have been formulated. Many models emphasize the role of climatic conditions on the dynamics of WNV, as temperature plays a critical role in the life-history traits of mosquitoes and virus replication. Thereby, temperature-dependent functions are used to model mosquito and virus development rates (Heidecke et al., 2025; Laperriere et al., 2011; Mbaoma et al., 2024). These are included in ordinary differential equations (ODEs) to analyze the dynamics of WNV. However, local transmission dynamics are generally modeled within the vector and host populations without considering the geographic spread. Fesce et al. (2022) estimated the basic reproductive number of the WNV infection and used it to quantify the infection spread in mosquitoes and birds, to classify areas with temperatures suitable for the outbreak of WNV. Other studies focus on the control of WNV, e.g., Bowman et al. (2005) simulated mosquito reduction methods and personal protective measures. A potential spill-over from birds to equids and humans was modeled by Laperriere et al. (2011) using temperature-driven parameters and providing values for parameters that govern the developments and transitions from one compartment to another. Other models analyzed differences in seasonal trends using sinusoidal (trigonometric) functions to represent seasonal variations of the WNV disease dynamics (Abdelrazec et al., 2015; Cruz-Pacheco et al., 2009; Moschini et al., 2017). Such models usually show seasons when mosquitoes are most likely to be active.
@@
Different modeling frameworks also include the role of migratory birds in the transmission of WNV. Bergsman et al. (2016) formulated an ODE model that investigates the interplay of WNV in both resident and migratory hosts. The model considers the yearly seasonal outbreaks that depend primarily on the number of susceptible migrant birds entering the local population each season, demonstrating that the early growth rates of seasonal outbreaks are more influenced by the migratory than the resident bird population. For Germany, just recently, Mbaoma et al. (2024) developed an ODE model for WNV in Germany by incorporating the complete life cycle of Cx. pipiens as a vector, resident birds, migratory birds, and humans as hosts. The study concluded that short-distance migratory birds such as the hooded crow were identified as important hosts in maintaining the enzootic transmission cycle of WNV in Germany.
@@
Although ODE-based models allow us to understand the transmission dynamics of WNV, they do not generally consider the spatial movement of hosts and vectors across different regions without incorporating additional types of models, like multi-patch models or distance dispersal kernels. For example, Bhowmick et al. (2023) used an ODE-based modeling approach in combination with different distance dispersal kernels to describe the spreading process of WNV through migratory birds in Germany. As a promising alternative, researchers have incorporated the effects of spatial variations and disease spread across different geographical regions in ODE-based models, using partial differential equation (PDE) models. Lin and Zhu (2017) developed a reaction-diffusion PDE model for the spatial spread of WNV in mosquitoes and birds in North America using free-boundary conditions. Another study of the WNV propagation was performed by Maidana and Yang (2009), where they proposed a spatial model for analyzing the WNV spread across the USA, calculating the traveling wave speed. The model considers a system of PDEs that describe the movement of mosquito and avian populations using the diffusion and advection processes. Results from this study show that mosquito movements do not play a major role in the spread, while bird advection is an important factor for lower mosquito biting rates.
@@
Maidana and Yang (2008, 2013) also contributed a lot to the spatial modeling of WNV using PDE models, and their framework is the basis of this study. Our model differs from previous WNV PDE models available in the literature, as we consider four different hosts, i.e., resident birds, migratory birds, humans, and equids, allowing us to capture the changes in the transmission patterns of WNV. Furthermore, we simulate the model in Germany, accounting for spatial heterogeneity, identify hotspots, and give insights into the contribution of migratory birds to WNV spread. In Germany, WNV was first detected in August 2018 in birds from the south-eastern region of Berlin (Ziegler et al., 2019). Since then, WNV has been established in Germany and has spread. Germany lies along migration routes of birds between Europe, the Mediterranean, and Africa (Musitelli et al., 2018). A study conducted by Nussbaumer et al. (2021) indicates that most migratory birds arrive in the northeastern region of Germany during spring and leave the country in early autumn via the southwestern direction. First WNV strains confirmed in Germany were closely related to strains previously detected in Austria and the Czech Republic (Ziegler et al., 2019), indicating a potential route of introduction, and their ability to overwinter in Cx. pipiens (Kampen et al., 2021) allows the long-term persistence of the virus.
@@ 2 Model formulation
Building on ODE-based models developed by Laperriere et al. (2011) and Mbaoma et al. (2024), we formulate a model for the spatial spread of WNV in Germany between 2019 and 2024. Our spatial model consists of a system of semi-linear parabolic PDEs governed by diffusion, advection, and reaction processes, together with temperature-dependent mosquito biting, latency, and mosquito mortality rates. In addition, our model includes the population of mosquitoes, resident and migratory birds. The population of mosquitoes at time and geographic location is divided into three compartments: susceptible exposed and infectious mosquitoes with the total population given by Susceptible mosquitoes get infected by biting an infectious resident bird and/or migratory bird with a temperature-dependent mosquito biting rate of
and a probability of acquiring the infection of and from resident and migratory birds respectively. The force of infection on mosquitoes is thus given by
where and are the environmental carrying capacities of the resident and migratory birds. A successful contact between infectious birds and susceptible mosquitoes at location and time sends susceptible mosquitoes to the exposed class. The latency rate on mosquitoes was fitted by Heidecke et al. (2025), and is given by
@@
defined by Laperriere et al. (2011). The host population consists of amplification hosts: resident birds and migratory birds. The population of hosts is divided into susceptible exposed , infected recovered and dead hosts, with and for easier readability, we denote resident birds by subscript and migratory birds by Laperriere et al. (2011) modeled the recruitment of susceptible birds using a gamma distribution, accounting for the observed seasonal birth rate of the American crow, which was the bird species observed to be mainly affected by the WNV in the Minneapolis metropolitan area. To reduce the non-linearity of our model, which has an impact on the numerical simulation of the PDE model, we compared our ODE simulation using a recruitment rate of but found no significant differences in the numerical results of our ODE version and theirs. Thus, our resident and migratory birds are recruited into the susceptible state at a rate of and respectively. Susceptible resident and migratory birds get infected by infectious mosquitoes, with a biting rate defined in Equation (1), and a probability of acquiring the infection of and respectively. The force of infection on resident birds is given by
@@
The intrinsic incubation period of resident and migratory birds is given by and respectively, while their recovery rates are given by and respectively. The proportions of resident birds and of migratory birds recover, while the proportions and die at the rates and respectively. The natural death rate of resident birds and migratory birds is given by with
@@ 2.1 Diffusion and advection processes
Diffusion refers to the process by which matter is transported from one part of a system to another as a result of random molecular motions (Crank, 1979). The concept of diffusion dates back to the 19th century, when Fick (1855) recognized the transfer of heat by conduction due to random molecular motions and quantified this process by adopting the mathematical equation of heat conduction previously derived by baron de Fourier (1822). To date, the diffusion process has gained significant attention, with mathematical biology being one of the application areas. In ecological modeling, the diffusion process is used to capture the random movement of species from regions of higher concentration to regions of lower concentration (Murray, 2007), which has an impact on their spatial distribution patterns. On the other hand, the advection process refers to the directional flow of a substance by a known velocity vector field (Boyd, 2001). Diffusion and advection processes combined with the ODE terms form an advection-diffusion-reaction PDE system that models how a phenomenon spreads in both time and space. Reaction terms are responsible for the local transmission at each location, while diffusion and advection are responsible for the spatial movement of species, including infectious specimens. In this study, we define diffusion for mosquitoes and both resident and migratory birds, as they locally move in search of resting/breeding sites, food, and other needs. The advection process is only defined for migratory birds, as we can account for their seasonal directional migration in the northeast direction during spring to late spring and fly out in the southwest direction starting from autumn to late autumn (Nussbaumer et al., 2021). The diffusion term is denoted by the Laplacian operator with diffusion coefficients for mosquitoes, resident, and migratory birds denoted by in respectively. Moreover, we choose the diffusion coefficients in a way that the diffusion of mosquitoes is less than that of resident and migratory birds, as mosquitoes are considered insects with a small radius of action (Harrington et al., 2005; Maidana and Yang, 2009; Verdonschot and Besse-Lototskaya, 2014). The diffusion coefficient of resident birds should also be less than that of migratory birds, to highlight the higher mobility of migratory birds. The advection term is denoted by the gradient operator with a velocity deg/day, in a two-dimensional plane.
The schematic diagram of the model under study is shown in Figure (1), while the system of equations describing the dynamics of the WNV model under study is shown in system (7).
The total populations of mosquitoes, resident birds, migratory birds, humans, and equids are given by while the initial conditions are given by
@@
We assume the system moves in a region with a smooth boundary according to Fick’s law (Fick, 1855), so that in the initial conditions, where and subject to the homogeneous Neumann boundary conditions
and is the unit outer normal to
Based on available WNV data in Germany, birds of the Accipitridae family were the most affected bird species in 2018. Therefore, as also done by Mbaoma et al. (2024), we used parameters associated with northern goshawks and raptors for the resident bird, both of which fall under the Accipitridae family. Initial conditions of infectious residential birds are estimated to be high in Halle (Saale), the main WNV hot-spot in 2018, followed by Berlin, and lastly few birds in Laage, Poing, and Plessa. More precisely, there are 2 infectious resident birds in Berlin, 1 in Plessa, Poing, and Laage, and 7 in Halle (Saale), as described by the end-of-year 2018 cases (Ziegler et al., 2020). For simplicity, we assume the same locations for initial conditions of infectious migratory birds. In the initial condition of susceptible hosts and mosquitoes, we assume a spatially homogeneous distribution, and we assign the total populations of each host uniformly across the entire domain. The total initial population of susceptible resident birds is assumed to be 100,000,000, while susceptible migratory birds are set at 500,000. The total mosquito population is estimated to be 20 million susceptible mosquitoes and 80,000 infectious mosquitoes, which agrees with an approximate minimum infection rate of 0.355 per 1000 mosquitoes estimated by Bernard et al. (2001). For numerical stability, we set initial dead resident birds and migratory birds to 1.
The variables are summarized in Table (1).
Variable Definition Value susceptible mosquitoes exposed mosquitoes 0 infectious mosquitoes 80,000 susceptible resident birds exposed resident birds 0 infectious resident birds 12 recovered resident birds 0 dead resident birds 1 susceptible migratory birds 500,000 exposed migratory birds 0 infectious migratory birds 12 recovered migratory birds 0 dead migratory birds 1
The model parameters are summarized in Table (2).
Parameter Definition Value Source natural mortality rate of resident birds 0.00034 (Laperriere et al., 2011) natural mortality rate of migratory birds 0.00023 (Mbaoma et al., 2024) latency rate on mosquitoes Eqn (3) (Heidecke et al., 2025) latency rate on resident birds 0.196 (Mbaoma et al., 2024) latency rate on migratory birds 0.285 (Mbaoma et al., 2024) removal rate of infectious resident birds 0.867 (Mbaoma et al., 2024) removal rate of infectious migratory birds 0.4 (Mbaoma et al., 2024) proportion of dead resident birds 0.655 (Mbaoma et al., 2024) proportion of dead migratory birds 0.103 (Mbaoma et al., 2024) probability of a new resident bird infection 0.9 (Mbaoma et al., 2024) probability of a new migratory bird infection 0.97 (Mbaoma et al., 2024) probability of a new mosquito infection from a resident bird 0.4 (Mbaoma et al., 2024) probability of a new mosquito infection from a migratory bird 0.7 (Mbaoma et al., 2024) mosquito biting rate Eqn (1) (Rubel et al., 2008) natural mortality rate of mosquitoes Eqn (4) (Tran et al., 2013) carrying capacity of mosquito population defined by the authors carrying capacity of resident bird population defined by the authors carrying capacity of migratory bird population defined by the authors diffusion coefficient of mosquitoes Table (3) inferred from the data diffusion coefficient of resident birds Table (3) inferred from the data diffusion coefficient of migratory birds Table (3) inferred from the data advection vector for migratory birds section (6) inferred from the data

### v2

@@ Abstract
West Nile virus (WNV) is a climate-sensitive mosquito-borne arbovirus circulating between mosquitoes of the genus Culex and birds, with a potential spillover to humans and other mammals. Recent trends in climatic change, characterized by early and/or prolonged summer seasons, increased temperatures, and above-average rainfall, probably facilitated the spread of WNV in Europe. In this work, we formulate a spatial WNV model consisting of a system of parabolic partial differential equations (PDEs), using the concept of diffusion and advection in combination with temperature-dependent parameters, i.e. mosquito biting rate, extrinsic incubation rate, and mortality rate. Diffusion represents the random movement of both mosquitoes and hosts across space, while advection captures the directed movement of migratory birds. The model is first studied mathematically, and we show that it has non-negative, unique, and bounded solutions in time and space. Numerical simulations of the PDE model are performed using temperature data for Germany (2019 - 2024). Results obtained from the simulation showed a high agreement with the reported WNV cases among birds and equids in Germany. The observed spreading patterns from the year 2018 to 2022 and the year 2024 were mainly driven by temperature in combination with diffusion processes of hosts and vectors. Only during the year 2023, the additional inclusion of advection for migratory birds was important to correctly predict new circulation areas in Germany.
@@ 1 Introduction
West Nile virus (WNV) is an arbovirus of the Japanese encephalitis antigenic complex of the Flaviviridae family. It was first detected in a febrile patient from the West Nile district of Northern Uganda in 1937 (Smithburn1940), but is now considered the most geographically widespread arbovirus in the world (Ciota2017). The virus circulates in an enzootic cycle between Culex mosquitoes and birds, with regular spillovers to humans, equids, and other mammals (Calisher1989; Hublek1999; Kramer2008). Birds are considered amplifying hosts as they develop sufficient levels of viremia, allowing for the infection of mosquitoes, while humans, equids, and other mammals are considered dead-end hosts (Rappole2003). Most humans infected with WNV are asymptomatic, but few may develop fever, headache, rash, among others (Petersen2013), including the risk of a fatal outcome (Campbell2002). Infected equids often exhibit ataxia, lethargy, and weakness as signs of the infection (Salazar2004). Multiple theories have been discussed regarding the introduction of new WNV cases in a previously unaffected region. There is evidence that migratory birds play an important role in the spread of the pathogen (Rappole2003). For example, birds captured during their seasonal migration have tested positive for WNV in Israel (ernek1977arboviruses; Malkinson2002; NIR1967; watson1972ectoparasite). Migratory birds likely spread the virus to stopover sites in Europe as they travel from the south to the north (hannoun1972role; Hublek1999). In contrast, considering the speed and duration of flights of migratory birds in combination with the duration of viremia, other researchers argued that the short-distance movement of non-migratory birds better explains the spread pattern of WNV (Rappole2003).
@@
Several mathematical models for WNV have been formulated. Many models emphasize the role of climatic conditions on the dynamics of WNV, as temperature plays a critical role in the life-history traits of mosquitoes and virus replication. Thereby, temperature-dependent functions are used to model mosquito and virus development rates (Heidecke2024; Laperriere2011; Mbaoma2024). These are included in ordinary differential equations (ODEs) to analyze the dynamics of WNV. However, local transmission dynamics are generally modeled within the vector and host populations without considering the geographic spread. Fesce2022 estimated the basic reproductive number of the WNV infection and used it to quantify the infection spread in mosquitoes and birds, to classify areas with temperatures suitable for the outbreak of WNV. Other studies focus on the control of WNV, e.g., BOWMAN2005 simulated mosquito reduction methods and personal protective measures. A potential spill-over from birds to equids and humans was modeled by Laperriere2011 using temperature-driven parameters and providing values for parameters that govern the developments and transitions from one compartment to another. Other models analyzed differences in seasonal trends using sinusoidal (trigonometric) functions to represent seasonal variations of the WNV disease dynamics (abdelrazec2015dynamics; CruzPacheco2009; Moschini2017). Such models usually only show seasons when mosquitoes are most likely to be active, without factoring in climatic effects.
@@
Different modeling frameworks also include the role of migratory birds in the transmission of WNV. DBergsman2016 formulated an ODE model that investigates the interplay of WNV in both resident and migratory hosts. The model considers the yearly seasonal outbreaks that depend primarily on the number of susceptible migrant birds entering the local population each season, demonstrating that the early growth rates of seasonal outbreaks are more influenced by the migratory than the resident bird population. For Germany, just recently, Mbaoma2024 developed an ODE model for WNV by incorporating the complete life cycle of Cx. pipiens as a vector; resident birds, migratory birds, and humans as hosts. The study concluded that short-distance migratory birds such as the hooded crow were identified as important hosts in maintaining the enzootic transmission cycle of WNV in Germany. These ecological patterns for bird migration provide the fundamental biological basis for incorporating migratory birds into our model. This is mainly because the bird movement patterns may determine where the virus can be introduced and how fast it can spread during the transmission seasons.
@@
Although ODE-based models allow us to understand the transmission dynamics of WNV, they do not generally consider the spatial movement of hosts and vectors across different regions without incorporating additional types of models, like multi-patch models or distance dispersal kernels. For example, Bhowmick2023 used an ODE-based modeling approach in combination with different distance dispersal kernels to describe the spreading process of WNV through migratory birds in Germany. As a promising alternative, researchers have incorporated the effects of spatial variations and disease spread across different geographical regions in ODE-based models, using partial differential equation (PDE) models. Lin2017 developed a reaction-diffusion PDE model for the spatial spread of WNV in mosquitoes and birds in North America using free-boundary conditions. Another study of the WNV propagation was performed by Maidana2009, where they proposed a spatial model for analyzing the WNV spread across the USA, calculating the traveling wave speed. The model considers a system of PDEs that describe the movement of mosquito and avian populations using the diffusion and advection processes. Results from this study show that mosquito movements do not play a major role in the spread, while bird advection is an important factor for lower mosquito biting rates.
@@
MAIDANA2008; MAIDANA2013 also contributed a lot to the spatial modeling of WNV using PDE models, and their framework is the basis of this study. Our model differs from previous WNV PDE models available in the literature, as we consider two different hosts, i.e., resident and migratory birds, allowing us to capture the changes in the transmission patterns of WNV. Furthermore, we simulate the model in Germany, accounting for spatial heterogeneity, identify hotspots, and give theoretical insights into the contribution of migratory birds to WNV spread. In Germany, WNV was first detected in August 2018 in birds from the Southeastern region of Berlin (Ziegler2019). Since then, WNV has been established in Germany and has spread. Germany lies along migration routes of birds between Europe, the Mediterranean, and Africa (Musitelli2018). A study conducted by Nussbaumer2021 indicates that most migratory birds arrive in the northeastern region of Germany during spring and leave the country in early autumn via the southwestern direction. First WNV strains confirmed in Germany were closely related to strains previously detected in Austria and the Czech Republic (Ziegler2019), indicating a potential route of introduction, and their ability to overwinter in Cx. pipiens (Kampen2021) allows the long-term persistence of the virus.
@@ 2 Model formulation
Building on ODE-based models developed by Laperriere2011 and Mbaoma2024, we formulate a model for the spatial spread of WNV in Germany between 2019 and 2024. Our spatial model consists of a system of semi-linear parabolic PDEs governed by diffusion, advection, and reaction processes, together with temperature-dependent mosquito biting, latency, and mortality rates. In addition, our model includes the population of mosquitoes, resident and migratory birds. The population of mosquitoes at time and geographic location is divided into three compartments: susceptible exposed and infectious with the total population given by Susceptible mosquitoes get infected by biting an infectious resident bird and/or migratory bird at a temperature-dependent mosquito biting rate
and a probability of acquiring the infection of and from resident and migratory birds respectively. The force of infection on mosquitoes is therefore given by
where and are the total resident and migratory bird populations. A successful contact between infectious birds and susceptible mosquitoes at location and time sends susceptible mosquitoes to the exposed class. The latency rate on mosquitoes was fitted by Heidecke2024, and is given by
@@
defined by Laperriere2011. Spatial and temporal plots of the thermal functions (equations 1, 2 and, 3) are shown in Appendix S1. The host population consists of amplification hosts: resident birds and migratory birds. The population of hosts is divided into susceptible exposed , infected recovered and dead hosts, with and for easier readability, we denote resident birds by subscript and migratory birds by Laperriere2011 modeled the recruitment of susceptible birds using a gamma distribution, accounting for the observed seasonal birth rate of the American crow, which was the bird species observed to be mainly affected by the WNV in the Minneapolis metropolitan area. To reduce the non-linearity of our model, which has an impact on the numerical solution of the PDE model, we compared our ODE simulation using a recruitment rate of but found no significant differences in the numerical results of our ODE version and theirs. Thus, our resident and migratory birds are recruited into the susceptible state at a rate of and respectively. Susceptible resident and migratory birds get infected by interacting with infectious mosquitoes, with a biting rate defined in Equation (1), and a probability of acquiring the infection of and respectively. In mosquito-borne disease models, the mosquito to bird ratio is very important as it determines the per bird biting pressure (Bhowmick2020; Laperriere2011; Mbaoma2024). Moreover, Vogels2017 argue that mosquito abundance is constrained by the availability of host blood meals, and bird density directly influences mosquito feeding success. As a result, the number of bites received by an individual bird depends primarily on the ratio of mosquitoes to birds, rather than on mosquito or bird density alone. Because of this, we introduce the mosquito to bird ratio parameters for residence and migratory birds, respectively. In WNV models, Bhowmick2020 simulated their ODE model with a mosquito to bird ratio ranging from 10-30, while Laperriere2011 used a ratio of 30. Another study by deFreitasCosta2024, which aimed at investigating the conditions necessary for the circulation of WNV in Germany and the Netherlands, found that larger proportions of susceptible birds and higher vector-host ratios were required. The study showed that WNV circulation in the Netherlands was rarely sustained when vector-host ratios fell below approximately 1:100.
The force of infection on resident birds is given by
@@
The latency rates on resident and migratory birds are given by rB and mB respectively, while their recovery rates are given by and Proportions of resident birds and of migratory birds recover from the infection, while the proportions rB and mB die at the rates and mB respectively. Resident and migratory bird species can also die naturally at rates rB and
@@ 2.1 Diffusion and advection processes
Diffusion refers to the process by which matter is transported from one part of a system to another as a result of random molecular motions (crank1979mathematics). The concept of diffusion dates back to the 19th century, when Fick1855 recognized the transfer of heat by conduction due to random molecular motions and quantified this process by adopting the mathematical equation of heat conduction previously derived by baron1822theorie. To date, the diffusion process has gained significant attention, with mathematical biology being one of the application areas. In ecological modeling, the diffusion process is used to capture the random movement of species from regions of higher concentration to regions of lower concentration (murray2007mathematical), which has an impact on their spatial distribution patterns. On the other hand, the advection process refers to the directional flow of a substance by a known velocity vector field (boyd2001chebyshev). Diffusion and advection processes combined with the ODE terms form an advection-diffusion-reaction PDE system that models how a phenomenon spreads in both time and space. Reaction terms are responsible for the local transmission at each location, while diffusion and advection are responsible for the spatial movement of species, including infectious specimens.
In this study, we define diffusion for mosquitoes and both resident and migratory birds, as they locally move in search of resting/breeding sites, food, and other needs. The advection process is defined for migratory birds only, as we can account for their seasonal directional migration patterns. In contrast, diffusion is denoted by the Laplacian operator with diffusion coefficients for mosquitoes, resident, and migratory birds denoted by in respectively. Moreover, we choose the diffusion coefficients in a way that the diffusion of mosquitoes is less than that of resident and migratory birds, as mosquitoes are considered insects with a small radius of action (HARRINGTON2005; Maidana2009; Verdonschot2014). A lower diffusion coefficient is assigned to resident birds compared to migratory birds to highlight the greater mobility of migratory birds. To improve interpretability of diffusion coefficients, we define the root mean squared displacement (RMSD). The RMSD describes the dispersal distance species move from the initial source under diffusion alone after a time (Einstein1905; metelmann2019development), and it is calculated using the formula:
where is the spatial dimension, is the diffusion coefficient and is the time step. We denote the advection term by the gradient operator with a velocity km/day, in a two-dimensional plane. The schematic diagram of our PDE model and the governing system of equations are shown in Figure (1) and system (6).
The total populations of mosquitoes, resident birds and migratory birds are given by while the initial conditions are given by:
@@
We assume the system moves in a region with a smooth boundary according to Fick’s law (Fick1855), so that in the initial conditions, where and subject to the homogeneous Neumann boundary conditions:
and is the unit outer normal to . The state variables and initial conditions are summarized in Table (S1) of the supplementary file, while model parameters are shown in Table (1).
Parameter Definition Value Source natural mortality rate of resident birds 0.0005 (Mbaoma2024) natural mortality rate of migratory birds 0.00023 (Mbaoma2024) latency rate on mosquitoes Eqn (2) (Heidecke2024) latency rate on resident birds 0.196 (Mbaoma2024) latency rate on migratory birds 0.285 (Mbaoma2024) removal rate of infectious resident birds 0.867 (Mbaoma2024) removal rate of infectious migratory birds 0.4 (Mbaoma2024) proportion of dead resident birds 0.655 (Mbaoma2024) proportion of dead migratory birds 0.103 (Mbaoma2024) probability of a new resident bird infection 0.97 (Mbaoma2024) probability of a new migratory bird infection 0.9 (Mbaoma2024) probability of a new mosquito infection from a resident bird 0.4 (Mbaoma2024) probability of a new mosquito infection from a migratory bird 0.7 (Mbaoma2024) mosquito biting rate Eqn (1) (rubel2008) natural mortality rate of mosquitoes Eqn (3) (Laperriere2011) mosquito to resident bird ratio 15 inferred from the data mosquito to migratory bird ratio 15 inferred from the data diffusion coefficient of mosquitoes Table (2) inferred from the data diffusion coefficient of resident birds Table (2) inferred from the data diffusion coefficient of migratory birds Table (2) inferred from the data advection vector for migratory birds [1;1] & [-1;-1] Estimated section (5)